• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dwang

    Log entry time 12:48:41 on October 5,2009

    Entry number 293271

    keyword=compton tuned

    compton repeated the same problem as we had yesterday. The symptom is that the high voltage keeps tripping and the background rate is high and not correlated with laser. Now we are pretty sure that the solution is to adjust the "phase of the laser" to minimize the background rate. Looks like this time we couldn't get back to the low background as before (100 Hz), but the background is less than 200, so it should be fine. After that, MCC did a vertical scan to optimize the rate. Now things looks back to normal. New production run started.

    FIGURE 1